OpenStudy (anonymous):

find f^-1(x) f(x)=(x-2)^1/3

OpenStudy (anonymous):

you can start by writing \[x=\sqrt[3]{y-2}\] and solve for \(y\)

OpenStudy (anonymous):

in two steps 1) cube both sides 2) add 2 to both sides

OpenStudy (anonymous):

y=x^3+2?

OpenStudy (anonymous):

\[y = \sqrt[3]{x-2} \Rightarrow y^3=x-2 or x = y^3+2 \]

OpenStudy (anonymous):

great thanx